# Theorem of the heart for Weibel’s homotopy -theory

### 09 Mar 2026 v1

@@ 4.1. Estimates for -groups
By Proposition 2.22 (ii) the functor induces an equivalence By Proposition 4.4 the functor takes to hence it takes to It follows by adjunction that the functor is fully faithful, and it induces isomorphisms on To see that it is essentially surjective, note that its essential image is closed under extensions and filtered colimits, and it contains for Therefore, this essential image coincides with ∎
@@
and denote by resp. the (strongly continuous) left adjoint to resp. right adjoint to Then the functors are isomorphic, and they are left adjoint to the forgetful functor. The functors and are left inverse to the functor
@@
We construct the -structures inductively, simultaneously proving the left -exactness assertion. We are already given the -structure on By assumption on and by Proposition 4.4, the functor is left -exact.
Suppose that for some we already have the claimed -structures on for such that for these the functor is left -exact. Applying Proposition 4.7, we obtain a unique compactly assembled continuously bounded -structure on such that the transition functor is -exact. We need to prove that the functor
@@ 4.2. Estimates for higher nil groups
We recall that for and for the higher nil spectrum is defined as where we take the reduced part in the simplicial sense and we consider the assignment as a simplicial -ring, see [E25c, Section 8]. On the level of homotopy groups we have
@@ 5. Theorem of the heart for
To prove this implication we use a general observation on the monads which are (non-deformed) tensor algebras.
@@
compatible with the transition maps. As in Proposition 4.6, we consider the functor right inverse to Recall that generates by construction, and by part (iii) of loc.cit., the adjunction counit is a split monomorphism, giving a direct sum decomposition Consider the morphism of monads corresponding by adjunction to the inclusion of as a direct summand. Put Then the maps of monads induce the chain of (strongly continuous exact) functors
@@
Applying to the functor we see that Hence, we have an isomorphism
@@ 7.1. Estimates for theorems of the heart and dévissage
There exist small -linear abelian categories such that the inclusion functor is exact, each object of is an extension of two objects of and the map is not surjective.
@@
Since the functors commute with filtered colimits, it suffices to consider the case when is equivalent to the category of finite projective right modules over a connective -ring such that is (left or right) artinian. Let and Then by [BGT13, Theorem 9.53] and (for example) by the proof of [LT19, Corollary 3.5] we have where is the Jacobson radical. The ring is a finite product of matrix algebras over skew-fields, hence we may assume that is a skew-field. In this case all the spectra vanish for ∎
@@
(i) Let be the cuspidal cubic curve over Denote by the category of vector bundles of finite rank on Denote by the dg category of acyclic bounded complexes of objects of By [Nee21, Lemma 1.2] this category has a bounded -structure whose heart is equivalent to the category of effaceable functors (see also Subsection 1.1, discussion before Proposition 1.3).
@@
We claim that for some the map is not an isomorphism, which would prove (i). It suffices to prove that the map is not an isomorphism.
By Proposition 2.25 (ii) we have so we need to show that this nil group is non-zero. By [Nee21, Proposition 2.4] the realization functor is an equivalence. Denoting by the category of with a split exact structure, we obtain a short exact sequence
@@
First we construct We use the construction from [E25b, Theorem 0.3] and recall the assertions of loc. cit. We denote by by the quasi-abelian category of pairs where and is a subobject such that both and are in The forgetful functor induces a quotient functor on bounded derived categories. We define to be its kernel, so we have a short exact sequence
@@
Both functors and are isomorphic to the derived functor of the inclusion Hence, by assumption the map is not surjective. In particular, the map is not a monomorphism, where is the right adjoint to But the functor is left inverse to the realization functor, hence the map is not surjective, as required.
Now suppose that for some we have constructed a -linear small -category with required properties such that the map is not surjective. We will construct the -category as follows. Put and consider the exact category with the induced exact structure. Consider another exact category from Proposition 3.7 with the same underlying category as (we use the small category version of loc. cit.). Recall from loc. cit. that a morphism in is an inclusion if both and are monomorphisms. We define via the short exact sequence
@@ 7.2. Estimates for higher nil groups of abelian categories
Moreover, we have
We elaborate on the meaning of the right hand side of (7.3). First, by we mean the negative cyclic homology over considered as a spectrum. The superscript “red” means the reduced part in the simplicial sense, where we consider the assignment as a functor see for example [E25c, Section 8].
@@
The fully faithful functor in (7.5) is the familiar categorical resolution of singularities from [KL15]. It is induced by the pullback functors: to the singular point and to the normalization
@@
Consider the universal finitary localizing invariant of -linear idempotent-complete stable categories We have an isomorphism This gives a natural commutative square in the category
@@
It follows easily from the [E25c, Theorem 9.3] that for any the spectrum is naturally a -module, where has cohomological degree (recall that ). Moreover, it follows from [E25c, Theorem 9.1] that is acting locally nilpotently on since by loc. cit. we have The proof of Theorem 7.3 shows that for the abelian category we have an isomorphism of -modules.
@@ 8.1. Sheaves on locally compact Hausdorff spaces
For simplicity we consider only the continuously bounded -structures, and we only consider the categories of sheaves with values in a fixed category (not in a presheaf of categories like [E24, Sections 6.2-6.4]).
@@ 8.4. Concluding remarks
We covered only some very basic non-trivial examples of continuously bounded -structures and coherently assembled abelian categories. We briefly mention the following directions.
@@
One can construct a dualizable -category starting from a coherently assembled exact category Namely, denote by the same category with an exact structure from Proposition 1.32. Then we have a short exact sequence in
One can show that the category has a natural compactly assembled continuously bounded -structure, induced by the natural -structure on If is locally coherent, then this -structure is compactly generated, and the induced -structure on coincides with the (bounded) -structure from Proposition 1.3. In general the heart is equivalent to the category of sequential limit-preserving additive functors such that for any for any and for any compact morphism with there exists an exact projection in such that

### 16 Mar 2026 v2

@@ 4.1. Estimates for -groups
By Proposition 2.22 (ii) the functor induces an equivalence By Proposition 4.4 the functor takes to hence it takes to It follows by adjunction that the functor is fully faithful, and it induces isomorphisms on To see that it is essentially surjective, note that its essential image is closed under extensions, and it contains for Therefore, this essential image coincides with and we have an equivalence of categories
It remains to observe that this equivalence automatically reflects exactness: in both categories the short exact sequences are precisely all fiber-cofiber sequences. ∎
@@
and denote by resp. the (strongly continuous) left adjoint to resp. right adjoint to Then the functors are isomorphic, and they are left adjoint to the forgetful functor. The functors and are left inverses to the functor
@@
We construct the -structures inductively, simultaneously proving the left -exactness assertion. We already have the -structure on By assumption on and by Proposition 4.4, the functor is left -exact.
Suppose that for some we have the claimed -structures on for such that for these the functor is left -exact. Applying Proposition 4.7, we obtain a unique compactly assembled continuously bounded -structure on such that the transition functor is -exact. We need to prove that the functor
@@ 4.2. Estimates for higher nil groups
We recall that for and for the higher nil spectrum is defined as where we take the reduced part in the simplicial sense and we consider the assignment as a simplicial -ring as explained in [E25c, Section 8]. On the level of homotopy groups we have
see [Wei89] for the -linear version.
@@ 5. Theorem of the heart for
To prove this implication we use a general observation on monads which are (non-deformed) tensor algebras.
@@
compatible with the transition maps. As in Proposition 4.6, we consider the functor right inverse to Recall that generates by construction, and by part (iii) of loc.cit., the adjunction counit is a split monomorphism, giving a direct sum decomposition Consider the morphism of monads corresponding by adjunction to the inclusion of as a direct summand. Put Then the maps of monads correspond to the chain of (strongly continuous exact) functors
@@
Applying to the functor we see that Hence, we have an equivalence
@@ 7.1. Estimates for theorems of the heart and dévissage
There exist small -linear abelian categories such that the (fully faithful) inclusion functor is exact, each object of is an extension of two objects of and the map is not surjective.
@@
Since the functors commute with filtered colimits, it suffices to consider the case when is equivalent to the category of finite projective right modules over a connective -ring such that is left or right artinian. Let and Then by [BGT13, Theorem 9.53] and by the nilinvariance of we have where is the Jacobson radical. The ring is a finite product of matrix algebras over skew-fields, hence we may assume that is a skew-field. In this case all the spectra vanish for ∎
@@
(i) Let be the cuspidal cubic curve over Denote by the category of vector bundles of finite rank on Denote by the dg category of acyclic bounded complexes of objects of By [Nee21, Lemma 1.2] this category has a bounded -structure whose heart is equivalent to the category of effaceable functors (this is a special case of Proposition 1.3).
@@
We claim that for some the map is not surjective, which would prove (i). It suffices to prove that the map is not surjective.
By Proposition 2.25 (ii) we have so we need to show that this nil group is non-zero. By [Nee21, Proposition 2.4] the realization functor is an equivalence. Denoting by the category with a split exact structure, we obtain a short exact sequence
@@
First we construct We use the construction from [E25b, Theorem 0.3] and recall the assertions of loc. cit. We denote by the quasi-abelian category of pairs where and is a subobject such that both and are in The forgetful functor induces a quotient functor on bounded derived categories. We define to be its kernel, so we have a short exact sequence
@@
Both functors and are isomorphic to the derived functor of the inclusion Hence, by assumption the map is not surjective. It follows that the map is not a monomorphism, where is the right adjoint to But the functor is left inverse to the realization functor, hence the map is not surjective, as required.
Now suppose that for some we have constructed a -linear small -category with required properties such that the map is not surjective. We will construct the -category as follows. Put and consider the exact category with the induced exact structure. Consider another exact category from Proposition 3.7 with the same underlying category as (we use the small category version of loc. cit.). Recall from loc. cit. that a morphism in is an inclusion if and only if both and are monomorphisms. We define via the short exact sequence
@@ 7.2. Estimates for higher nil groups of abelian categories
We have
We elaborate on the meaning of the right hand side of (7.3). First, by we mean the negative cyclic homology over considered as a spectrum [Tsy83, FT85, Con94]. The superscript “red” means the reduced part in the simplicial sense, where we consider the assignment as a functor see for example [E25c, Section 8].
@@
The fully faithful functor in (7.5) is the familiar categorical resolution of singularities from [KL15]. It is induced by the pullback functors: to the singular point and to the normalization The essential image of is in fact the strict pullback.
@@
Consider the universal finitary localizing invariant of -linear idempotent-complete stable categories Then we have an isomorphism This gives a natural commutative square in the category
@@
It follows directly from the [E25c, Theorem 9.3] that for any the spectrum is naturally a -module, where has cohomological degree (recall that ). Moreover, it follows from [E25c, Theorem 9.1] that is acting locally nilpotently on since by loc. cit. we have The proof of Theorem 7.3 shows that for the abelian category we have an isomorphism of -modules.
@@ 8.1. Sheaves on locally compact Hausdorff spaces
For simplicity we consider only the continuously bounded -structures, and we only consider the categories of sheaves with values in a fixed category (not in a presheaf of categories like in [E24, Sections 6.2-6.4]).
@@ 8.4. Concluding remarks
We covered only some very basic non-trivial examples of compactly assembled -structures and coherently assembled abelian categories. We briefly mention the following directions.
@@
One can construct a dualizable -category starting from a coherently assembled exact category Namely, (slightly abusing the notation) denote by the same category with an exact structure from Proposition 1.32. Then we have a short exact sequence in
One can show that the category has a natural compactly assembled continuously bounded -structure, induced by the natural -structure on If is locally coherent, then and the induced -structure on coincides with the (bounded) -structure from Proposition 1.3. In general the heart is equivalent to the category of sequential limit-preserving additive functors such that for any for any and for any compact morphism with there exists an exact projection in such that